OpenStudy (anonymous):

I WILL GIVE MEADEL AND FAN!!!!!!!!!!! Yesterday, a parking log lot had 10 cars parked cars parked in it. of cars parked in the parked in the parking lot, 40% were silver, 3 were red, and the rest were blue how many cars were blue? A.2 B.5 C.3 D.8

OpenStudy (anonymous):

C. 0.4*10=4 4+3=7 10-7=3

OpenStudy (anonymous):

thanks music life

OpenStudy (anonymous):

You're welcome! Anytime! @oregon234